For the structures shown below, state the number of pi electrons present in the molecule.

Is the molecule aromatic according to the Huckel c
Answer #1

Solution :-

In the first structure we have 3 rigns

And it have 6 triple bonds

So pi electrons = 6*2 = 12

But for the 3 ring we need (4*3)+2 = 14 electrons so that it become aromatic

So the structure 1 is not aromatic

Because it do not matches with Huckel criteria.

In the structure 2

We have 1 ring which has 3 double bonds so the total pi electrons in the ring are 3*2 = 6

So it obeys the Huckel rule so it is aromatic

Structure 2 is aromatic    because it matches with Huckel criteria.